can someone please help.
solve each equation for x
1. 4^(x + 1) = 4^(3x - 5)
2.8^2x = 8^(3x - 1)
3.10^6x = 100^(x + 6)

use your calculator to solve for x.round to the nearest thousandth
1.10^2.3 = x
2.e^x = 9
3.10^x = 15
4.e^ -1.5 = x
thanks for the help.I really appreciate it.

Answer by stanbon(48535)

1. 4^(x + 1) = 4^(3x - 5)
x+1 = 3x-5
2x = 6
x = 3
---------------------------
2.8^2x = 8^(3x - 1)
2x = 3x-1
x = 1
==================
3.10^6x = 100^(x + 6)
10^6x = 10^(2x+12)
6x = 2x+12
4x = 12
x = 3
-----------------------
use your calculator to solve for x.round to the nearest thousandth
1.10^2.3 = x
x = 199.526..
----------------------
2.e^x = 9
x = ln9
x = 2.1972...
------------------------
3.10^x = 15
x = log(15)
x = 1.17609...
--------------------------
4.e^ -1.5 = x
x = 0.22313...
=============================
Cheers,
Stan H.
Reply to stanbon@comcast.net or algebra.com